40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Oct 18 22:05:40 2009, kiwamu, Update, LSC, LSC timing issue rough.pngdetail.png
    Reply  Sun Oct 18 23:02:03 2009, Koji, Update, LSC, LSC timing issue 
       Reply  Mon Oct 19 10:00:52 2009, kiwamu, Update, LSC, RE: LSC timing issue 
Message ID: 2111     Entry time: Sun Oct 18 22:05:40 2009     Reply to this: 2113
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: LSC timing issue 

Today I made a measurement to research the LSC timitng issue as mentioned on Oct.16th.

*method

I put the triangular-wave into the OMC side (OMC-LSC_DRIVER_EXT) by AWG,

then looked at the transferred same signal at the LSC side (LSC_DARM_IN1) by using tdsdata.

I have compared these two signals in time domain to check whether they are the same or not.

In the ideal case it is expected that they are exactly the same.

 

*preliminary result

The measured data are shown in attached fig.1 and 2.

In the fig.1 it looks like they are the same signal.

However in fig.2 which is just magnified plot of fig.1, it shows a time-delay apparently between them.

The delay time is roughly ~50 micro sec.

The surprising is that the LSC signal is going beyond the OMC signal, although the OMC signal drives the LSC !!

We can say it is "negative delay"...

Anyway we can guess that the time stamp or something is wrong.

 

*next plan

Tomorrow I'm going to measure the transfer-function between them to see the delay more clearly.

( And I would like to fix the delay. )

Attachment 1: rough.png  62 kB  | Hide | Hide all
rough.png
Attachment 2: detail.png  45 kB  | Hide | Hide all
detail.png
ELOG V3.1.3-